Suggested Searches

This content has been archived.

ISS On-Orbit Status Report

    ISS Daily Summary Report – 2/17/2017

    NanoRack Cubesat Deployer (NRCSD) #10 Installation: This morning the crew installed two Quad deployers on the attachment mechanisms on the JEM Airlock (JEMAL) Slide Table. Only one Quad has satellites loaded, the second Quad is required for proper fit of the Multi-Layered Insulation (MLI) protective cover.  The deployment is scheduled for after the berthing and …

    Read Full Post

    ISS Daily Summary Report – 2/16/2017

    Transition to (X2) R15 Software:  As part of the R15 software transition, the Starboard 1 (S1) and Port 1 (P1) Multiplexer/Demultiplexers (MDMs) were transitioned to updated software S1P1 R6. Updates to the S1 and P1 MDMs include:  Longeron Fault Detection Isolation & Recovery (FDIR) updates to protect for Beta Gimbal Assembly (BGA) stall conditions Thermal …

    Read Full Post

    ISS Daily Summary Report – 2/15/2017

    Transition to (X2) R15 Software:  As part of the R15 software transition, the Payload (PL) Multiplexer/Demultiplexers (MDMs) were transitioned to updated software Payload Executive Processor R12 (PEP R12) today. Updates to the PL MDMs include:  Redundancy Management Time Reduction Additionally, the Lab-1 (LA-1) MDM was transitioned to Lab System-R5 (LSYS1-R5) software. Updates to the LA-1 …

    Read Full Post

    ISS Daily Summary Report – 2/14/2017

    Transition to (X2) R15 Software:  As part of the R15 software transition, two of the three Command and Control (C&C) Multiplexer/Demultiplexers (MDMs) were transitioned to updated software today. Updates to the C&C MDMs include:  Visiting Vehicle Support Implement modified capability for loading patches Modification to Rapid Depress Auto-Response Dual Integrated Communications Unit (ICU) Capability Increase …

    Read Full Post

    ISS Daily Summary Report – 2/13/2017

    Electrostatic Levitation Furnace (ELF) Sample Clean: This morning the crew disconnected the Sample Cartridges from within the ELF to exchange the samples and retrieve lost samples within the sample holders.  The crew reinserted the cartridge and reconnected the hoses and cables.  Ground teams will resume checkout and operations later this week. The ELF is an …

    Read Full Post

    ISS Daily Summary Report – 2/10/2017

    Japanese Experiment Module (JEM) ORU Transfer Interface (JOTI) and Robotics External Leak Locator (RELL) Installation on JEM Airlock (JEMAL) Slide Table: The crew installed the JOTI on the JEM A/L Slide Table. Afterwards, the crew installed the Robotics External Leak Locator (RELL) onto the JOTI, and retracted the Slide Table inside the JEMAL.  RELL operations …

    Read Full Post

    ISS Daily Summary Report – 2/09/2017

    Galley Rack Outfitting: Today the crew relocated the Node 1 16-Port Ethernet Switch from Node 1 Port-Deck location to the Galley Rack. All pre-existing cables remained mated to the switch during relocation, and the switch power feed will remain the same. Later today they will connect the two Microgravity Experiment Research Locker/Incubator (MERLIN) Ethernet connections …

    Read Full Post

    ISS Daily Summary Report – 2/08/2017

    Galley Rack Outfitting: The crew continued with Galley Rack Outfitting today. They mated power, data, and fluid umbilicals to the Galley Rack in preparation for activation. They also relocated the Food Warmer from the Node 1 Starboard location to the Node 1 Galley Rack. Following this relocation the Food Warmer will be powered from the …

    Read Full Post

    ISS Daily Summary Report – 2/07/2017

    Japanese Experiment Module (JEM) Exposed Facility (EF) High Definition Television (HDTV) Camera Installation: Today, the crew installed the HDTV on the EFU Adapter which was then installed on the JEM Airlock Slide Table. The Slide Table was then retracted into the JEM Airlock, and the JEM Airlock was depressurized and leak checked in preparation for operations …

    Read Full Post

    ISS Daily Summary Report – 2/06/2017

    H-II Transfer Vehicle (HTV)-6 Deorbit:  HTV-6 successfully completed all three deorbit burns over the weekend.  Loss of telemetry occurred Sunday 05-February at 15:11:21 GMT at an altitude of 80 km over the South Pacific near 50° S, 142° W.  Lighting Effects Vision Test: This morning the crew unstowed the Visual Performance Test hardware, then configured …

    Read Full Post

Subscribe to this blog

Show Past Archives